The sum of the ages of two brothers is 38 years. Four years ago, the age of the elder brother was the square of the younger brother's age.Find their ages.
Found 2 solutions by stanbon, ankor@dixie-net.com:
Answer by stanbon(75887)

The sum of the ages of two brothers is 38 years. Four years ago, the age of the elder brother was the square of the younger brother's age.Find their ages.
------
Equations:
a + b = 38 years
a-4 = (b-4)^2
--------------------
Rearrange:
a = 38-b
Substitute for "a" and solve for "b":
(38-b)-4 = b^2 - 8b + 16
b^2 - 7b -18 = 0
----
(b-9)(b+2) = 0
----
Positive solution:
b = 9 (one of the brother's age)
Solve for "a":
a + b = 38
a + 9 = 38
a = 29 (age of the other brother)
==================
Cheers,
Stan H.
==================

Answer by ankor@dixie-net.com(22740)